# On the Quest for Effectiveness in Human Oversight: Interdisciplinary Perspectives

### v1

@@ Abstract.
Human oversight is currently discussed as a potential safeguard to counter some of the negative aspects of high-risk AI applications. This prompts a critical examination of the role and conditions necessary for what is prominently termed effective or meaningful human oversight of these systems. This paper investigates effective human oversight by synthesizing insights from psychological, legal, philosophical, and technical domains. Based on the claim that the main objective of human oversight is risk mitigation, we propose a viable understanding of effectiveness in human oversight: for human oversight to be effective, the human overseer has to have (a) sufficient causal power with regards to the system and its effects, (b) suitable epistemic access to relevant aspects of the situation, (c) self-control over their own actions, and (d) fitting intentions for their role. Furthermore, we argue that this is equivalent to saying that a human overseer is effective if and only if they are morally responsible and have fitting intentions. Against this backdrop, we suggest facilitators and inhibitors of effectiveness in human oversight when striving for practical applicability. We discuss factors in three domains, namely, the technical design of the system, individual factors of human overseers, and the environmental circumstances in which the overseer operates. Finally, this paper scrutinizes the upcoming AI Act of the European Union – in particular Article 14 on Human Oversight – as an exemplary regulatory framework in which we study the practicality of our understanding of effective human oversight. By analyzing the provisions and implications of the European AI Act proposal, we pinpoint in how far that proposal aligns with our analyses regarding effective human oversight as well as how it might get enriched by our conceptual understanding of effectiveness in human oversight.
@@
We want to thank Sven Hetmank and Franz Lehr for their contributions to previous work that this paper is based on.
@@ 1. Introduction
Effective or meaningful [^1] Henceforth, we focus exclusively on the term “effective” and omit the term “meaningful”. Given their synonymous use in the context of human oversight, any statements regarding effectiveness also extend to meaningfulness. Furthermore we only are interested in human overseers, so we occasionally drop the first word in “human oversight” for the sake of simpler wording. human oversight is a notorious centerpiece of ethical guidelines (Jobin et al., 2019) and global legislation (European Parliament, 2024; Green, 2022) intended to govern the deployment of AI-based systems, particularly in high-risk contexts. The regulatory significance of human involvement in automated decision-making is unmistakable, as evidenced by a range of legislation (Green, 2022; Enqvist, 2023), including the overarching principles embedded in the General Data Protection Regulation of the EU (European Union, 2016) and more targeted legislation, such as the European AI Act (European Parliament, 2024), which mandates human oversight for high-risk AI applications. Research has made strides in defining and assigning responsibility for human oversight (Enqvist, 2023), or placing it within a broader risk management framework (Green, 2022). Human oversight has also been subject to criticism, with concerns being raised about the efficacy of oversight in practice, and its potential role as a mere vehicle to legitimize imperfect AI systems (Green, 2022; Zerilli et al., 2019).
Effective oversight has been discussed by many (Green, 2022; Enqvist, 2023; Laux, 2023; Methnani et al., 2021), however, a question that is crucial for this discussion has not been sufficiently addressed: When is human oversight effective?. The pressing need for conceptual clarity in this regard is evident, particularly given that the European Unions AI Act is to be implemented in 2024. Nevertheless, the definition and conditions for achieving effective human involvement remain underexplored, highlighting a critical gap in the current state of art of AI governance.
@@
  • We suggest a viable understanding of effectiveness in the context of human oversight based on the idea that the central objective of oversight is risk-mitigation. Namely, we propose that an overseer is effective if and only if they have (a) sufficient causal power with regards to the system and its effects, (b) suitable epistemic access to relevant aspects of the situation, (c) self-control over their own actions, and (d) fitting intentions for their role. We point out that (a)–(c) are jointly sufficient for moral responsibility.
@@
  • We exemplify facilitators and inhibitors of effectiveness in human oversight in three categories: the technical design of the system, the individual factors of the human overseer, and the environmental circumstances in which the overseer operates.
  • As a litmus test, we look at the details of the European AI Act, relative to our proposed conceptualization. We identify possible benefits and shortcomings of the AI Act. In particular, we argue that our conceptualization provides both, a more general and a more practically useful conceptualization of effective human oversight compared to the respective stipulations in the AI Act.
@@ 2. Related Work
Researchers have spelled out who will be in charge of (ensuring) human oversight and have analyzed how human oversight fits into a more global risk management concept (as, e.g., in the AI Act (European Parliament, 2024)) together with other approaches on accurate and robust systems, data-privacy adhering systems, and transparent systems (Enqvist, 2023; Laux, 2023). The concept of human oversight, however, also faces substantial criticism, as empirical evidence suggests that humans may be ineffective in supervising AI-based systems (Green, 2022). This has prompted a call for a more institutional oversight approach, emphasizing the need for empirical evidence regarding humans’ capability to oversee AI systems prior to enforcing the oversight (Green, 2022). This section will explore literature on effective oversight from different disciplines in more detail.
@@
Oversight is an essential aspect of various ethical guidelines and legislation on the use of AI-based systems in high-risk contexts. After the analysis of 41 policy documents around the globe, Green (Green, 2022) observes a call for limiting solely automated decisions (e.g., Article 22 of the European General Data Protection Regulation (European Union, 2016)), and an emphasis on the need for human discretion in AI-based decision-making (e.g., Canadian Directive on Automated Decision-Making (Government of Canada, 2023)). The requirement for effective oversight is particularly highlighted in the European AI Act. At the time of writing, a final version of the AI Act was approved by the EU Parliament in March 2024 (European Parliament, 2024) but has not yet been formally adopted. However, we will already use this final version, but will refer to earlier versions when relevant.
@@
The focus on human oversight seems to stem from the background of ensuring a human-centric development and deployment of AI-based systems. One of the most prominent concerns regarding the wide-range deployment of AI-based systems seems to be “that technological development and rationalized efficiency will take place at the cost of human agency and safety or rights” (Enqvist, 2023). An obvious way to counter such a techno-centric deployment of AI would be a human-centric development and deployment of AI-based systems. As Enqvist (Enqvist, 2023) notes: “Human centricity, as it has come to be (broadly) understood, does not only reflect that human needs are to be met by new technologies, but also incorporates the aim to safeguard individual rights and increase human well-being.” Indeed, in computer science a popular area of research is concerned with human-centered AI (HCAI) which investigates efficient collaboration between humans and AI systems (Bansal et al., 2019; Buçinca et al., 2021; Ehsan et al., 2022; Lai et al., 2023; Schoeffer et al., 2023; Tahaei et al., 2023).
@@
One of the foundational assumptions for the effectiveness of oversight measures is that human oversight should not be an end in itself (Laux, 2023), but that the involvement of humans in contexts where AI-based systems affect high-risk decision-making can make things better in some regard. For example, one hope is that humans may be better than AI-based systems at incorporating ethical considerations and social norms into decision-making contexts (Enqvist, 2023). Another hope is that humans might be better at judging single or unusual cases, for instance, in decisions that would affect individual human beings (Longoni et al., 2019). Additionally, associating the term “effective” with human oversight shows the hope that a human can effectively contribute to a joint decision situation together with an AI-based system, thus enabling decisions better than those of either the human or the system alone (Bansal et al., 2021; Buçinca et al., 2021; Schoeffer et al., 2023). Other hopes include increased safety (Koulu, 2020), improved accuracy (Jones, 2017), or better trustworthiness and more actual trust (Methnani et al., 2021). The most important objective of human oversight recently seems to be the mitigation of risks to fundamental rights, both directly (as one reading of the AI Act (European Parliament, 2024) suggests), or indirectly via, e.g., the goal to improve safety or the avoidance of ethically and socially undesirable outcomes (Enqvist, 2023).
@@
However, these foundational assumptions have received substantial criticism calling into question the usefulness of human oversight (Green, 2022; Zerilli et al., 2019). Critics (in particular (Green, 2022)) point to a lack of empirical evidence showing that we, as of now, can reliably integrate human and system abilities in a way that leads to better decisions compared to decisions that each single entity would produce in isolation (Bartlett and McCarley, 2020; Bansal et al., 2021; Rieger and Manzey, 2022). For instance, there seem to be substantial challenges with respect to achieving adequate trust in AI-based systems (Bansal et al., 2019; Lee and See, 2004; Hoff and Bashir, 2015; Zhang et al., 2020). Humans may overly rely on the outputs produced by automated systems, leading to situations where they do not detect erroneous or unfair outputs (Parasuraman and Manzey, 2010). In other situations, humans may have too little trust in AI-based systems or too high confidence in their own abilities which leads to people overriding actually accurate or fair system outputs (Green and Chen, 2019). Especially in high reliability contexts, automation bias (cf. Section 4.2) is a likely-to-observe phenomenon where people tend to use the outcome of an automated decision entity “as a heuristic replacement for vigilant information seeking and processing” ((Mosier et al., 1996) p. 205). This makes it particularly challenging for people to detect and adequately address system failures or erroneous outputs. Whereas research has identified several avenues to improve human abilities to oversee systems (e.g., training of human overseers (Bahner et al., 2008) augmenting system outputs with confidence scores (Bahner et al., 2008; Bansal et al., 2019)), the positive effects associated with these avenues seem to be inconsistent (Green, 2022). The criticism regarding human oversight could be expressed in an even more foundational way: The driving momentum behind many AI-based systems is the intention of producing decisions and actions that are, in some regard, better than what human experts could achieve (Green, 2022; Zerilli et al., 2019). Ironically, the very human beings whose imperfections are meant to be overcome by such an AI-based system end up in the role of overseers of the system (Zerilli et al., 2019). This is, for instance, a typical situation in certain areas of AI-based medical diagnostics (Haggenmüller et al., 2021; Nishida et al., 2022).
@@
However, it is also possible to draw more positive conclusions from the literature. There are relevant examples where human operators successfully identified errors made by automated systems (McBride et al., 2013; De-Arteaga et al., 2020). In addition, research has enhanced our understanding of factors that affect human abilities to detect errors made by automated systems (De-Arteaga et al., 2020; McBride et al., 2013; Lai et al., 2023; Schoeffer et al., 2023). Furthermore, ongoing research contributes to successfully unveiling design options to improve human-system collaboration (Lai et al., 2023; Vasconcelos et al., 2023). For example, recent work proposes to combine the strengths of machine and human decision making by using so-called conformal predictions for multi-label classifiers. There, the machine proposes a set of outputs (instead of a single output), which is guaranteed to contain the correct output with a probability of at least some pre-defined value (Straitouri et al., 2023). If this set contains only a single output, confidence is considered high enough to let the decision happen fully autonomously. If instead it contains more than one output, the choice is delegated to a human. It was shown that for certain tasks, the human-system decision accuracy indeed is higher when the human choice is restricted to the outputs that are in the proposed set (Straitouri and Rodriguez, 2023).
Overall, research reports mixed results regarding human oversight. However, the fact that effective human oversight is a popular requirement in legislation emphasizes the necessity of obtaining clarity on the very concept of effective human oversight. This clarity is crucial to better devise powerful laws and guidelines and for ensuring the successful and safe implementation of AI in high-risk contexts. Thus, in this paper we aim to derive high-level conditions for effectiveness in human oversight.
@@ 3. What is Effective Human Oversight?
We understand human oversight to be the supervision of a system by at least one natural person, typically with the authority to influence its operations or effects. [^2] While this understanding of human oversight will serve as our working definition, it is not the only possible candidate (cf. Section 6). For further analysis of the concept of human oversight, also see (Laux, 2023). This influence operates at various levels throughout the system’s lifecycle, including intervention during execution, reversal of faulty decisions, or adjustments of parameters to improve results (Methnani et al., 2021). Although the overseer may be assigned additional responsibilities, such as monitoring system performance or making engineering decisions, these tasks are not themselves part of human oversight. Even when having fixated a meaning for “human oversight”, it is not clear what “effective” means in this regard.
3.1. An objective-first approach to effectiveness of human oversight
@@
We want oversight to be useful. Therefore, it makes sense to define the effectiveness of human oversight in terms of how well it achieves its objectives. As we have discussed above, risk mitigation can be taken to be one of the main objectives of human oversight. This is the objective we will be focusing on. Additionally, effective human oversight not only mitigates risks but arguably may also yield other positive effects, such as improved incorporation of ethical considerations (Enqvist, 2023), or enhanced judgment in cases impacting individual well-being (Longoni et al., 2019).
@@
It is important to note, though, that the understanding of effectiveness we are about to suggest is not the only viable one. While offering a structured framework, it is just one among potentially numerous conceptualizations. One way to change our conceptualization of effectiveness in human oversight would be to put objectives other than risk mitigation into focus, such as improved accuracy, trustworthiness, actual trust, human autonomy, accountability and responsibility, or liability. Discussing them is beyond the scope of this paper, however, prima facie, there appears to be no inherent conflict between our proposal and these objectives. Furthermore, we believe that our conceptualization will be helpful beyond risk mitigation, since it will also be able to address other objectives of human oversight, including (moral) responsibility as we will demonstrate in Section 3.2.2.
@@
In line with the proposition that the main objective of effective human oversight is to mitigate risk, the main question therefore is: When does human oversight facilitate the mitigation of risks? We propose that this is the case if the human overseer meets the following four key conditions: causal power, epistemic access, self-control, and fitting intentions. We believe that this makes intuitive sense: Usually, risks can be reliably mitigated by a human only if there is an action available to them that averts the risk (causal power), they can actually perform that action (self-control), they know that the risk is imminent and which action could mitigate it (epistemic access), and they also want to mitigate it (fitting intentions). This intuition can be put more precisely, as is done in the remainder of this subsection. Additionally, Figure 1 displays the four conditions of effectiveness.
@@
Figure 1. The four conditions of effectiveness in human oversight and their relation to moral responsibility. The person in the diagram depicts the human overseer and the graph a schematic depiction of their current decision situation: the node on the left is their current position, the edges are the actions that they could perform, and the nodes on the right are the consequences of their actions.

### v2

@@ Abstract.
Human oversight is currently discussed as a potential safeguard to counter some of the negative aspects of high-risk AI applications. This prompts a critical examination of the role and conditions necessary for what is prominently termed effective or meaningful human oversight of these systems. This paper investigates effective human oversight by synthesizing insights from psychological, legal, philosophical, and technical domains. Based on the claim that the main objective of human oversight is risk mitigation, we propose a viable understanding of effectiveness in human oversight: for human oversight to be effective, the oversight person has to have (a) sufficient causal power with regard to the system and its effects, (b) suitable epistemic access to relevant aspects of the situation, (c) self-control, and (d) fitting intentions for their role. Furthermore, we argue that this is equivalent to saying that an oversight person is effective if and only if they are morally responsible and have fitting intentions. Against this backdrop, we suggest facilitators and inhibitors of effectiveness in human oversight when striving for practical applicability. We discuss factors in three domains, namely, the technical design of the system, individual factors of oversight persons, and the environmental circumstances in which they operate. Finally, this paper scrutinizes the upcoming AI Act of the European Union – in particular Article 14 on Human Oversight – as an exemplary regulatory framework in which we study the practicality of our understanding of effective human oversight. By analyzing the provisions and implications of the European AI Act proposal, we pinpoint how far that proposal aligns with our analyses regarding effective human oversight as well as how it might get enriched by our conceptual understanding of effectiveness in human oversight.
Human Oversight, AI Act, High-Risk AI, Law, Psychology
@@
We want to thank Sven Hetmank and Franz Lehr for their contributions to previous work that this paper is based on. Also, we want to thank three anonymous reviewers for their valuable feedback.
journalyear: 2024copyright: rightsretainedconference: ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3659051isbn: 979-8-4007-0450-5/24/06ccs: Social and professional topics Governmental regulationsccs: Social and professional topics Computing / technology policyccs: Human-centered computingccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligence
@@ 1. Introduction
Effective or meaningful [^1] Henceforth, we focus exclusively on the term “effective” and omit the term “meaningful”. Given their synonymous use in the context of human oversight, any statements regarding effectiveness also extend to meaningfulness. Furthermore, we are only interested in human oversight, so we occasionally drop the “human” and talk about “oversight” only for the sake of simpler wording. In response to a very insightful reviewer comment, we have also avoided the phrase “human overseer” because of its problematic connotation in US-American contexts. human oversight is a notorious centerpiece of ethical guidelines (Jobin et al., 2019) and global legislation (European Parliament, 2024a; Green, 2022) intended to govern the deployment of AI-based systems, particularly in high-risk contexts. The regulatory significance of human involvement in automated decision-making is unmistakable, as evidenced by a range of legislation (Green, 2022; Enqvist, 2023), including the overarching principles embedded in the General Data Protection Regulation of the EU (European Union, 2016) and more targeted legislation, such as the European AI Act (European Parliament, 2024a), which mandates human oversight for high-risk AI applications. Research has made strides in defining and assigning responsibility for human oversight (Enqvist, 2023), or placing it within a broader risk management framework (Green, 2022). Human oversight has also been subject to criticism, with concerns being raised about the efficacy of oversight in practice and its potential role as a mere vehicle to legitimize imperfect AI systems (Green, 2022; Zerilli et al., 2019).
Effective oversight has been discussed by many (Green, 2022; Enqvist, 2023; Laux, 2023; Methnani et al., 2021); however, a question that is crucial for this discussion has not been sufficiently addressed: When is human oversight effective? The pressing need for conceptual clarity in this regard is evident, particularly given that the European Union’s AI Act is due to be adopted in 2024. Nevertheless, the definition and conditions for achieving effective human involvement remain underexplored, highlighting a critical gap in the current state of the art of AI governance.
@@
  • We suggest a viable understanding of effectiveness in the context of human oversight based on the idea that the central objective of oversight is risk mitigation. Namely, we propose that an oversight person is effective if and only if they have (a) sufficient causal power with regards to the system and its effects, (b) suitable epistemic access to relevant aspects of the situation, (c) self-control over their own actions, and (d) fitting intentions for their role. We point out that (a)–(c) are jointly sufficient for moral responsibility.
@@
  • We exemplify facilitators and inhibitors of effectiveness in human oversight in three categories: the technical design of the system, the individual factors of the human who is in charge of oversight, and the environmental circumstances in which they operate.
  • As a litmus test, we look at the details of the European AI Act relative to our proposed conceptualization. We identify possible benefits and shortcomings of the AI Act. In particular, we argue that our conceptualization provides both a more general and a more practically useful conceptualization of effective human oversight compared to the respective stipulations in the AI Act.
@@ 2. Related Work
Researchers have spelled out who will be in charge of (ensuring) human oversight and have analyzed how human oversight fits into a more global risk management concept (as, e.g., in the AI Act (European Parliament, 2024a)) together with other approaches on accurate and robust systems, data-privacy adhering systems, and transparent systems (Enqvist, 2023; Laux, 2023). The concept of human oversight, however, also faces substantial criticism, as empirical evidence suggests that humans may be ineffective in supervising AI-based systems (Green, 2022). This has prompted a call for a more institutional oversight approach, emphasizing the need for empirical evidence regarding humans’ capability to oversee AI systems prior to enforcing the oversight (Green, 2022). This section will explore the literature on effective oversight from different disciplines in more detail.
@@
Oversight is an essential aspect of various ethical guidelines and legislation on the use of AI-based systems in high-risk contexts. After the analysis of 41 policy documents around the globe, Green (Green, 2022) observes a call for limiting solely automated decisions (e.g., Article 22 of the European General Data Protection Regulation (European Union, 2016)), and an emphasis on the need for human discretion in AI-based decision-making (e.g., Canadian Directive on Automated Decision-Making (Government of Canada, 2023)). The requirement for effective oversight is particularly highlighted in the European AI Act. At the time of writing, the AI Act has not yet been formally adopted. However, a final version was approved by the EU Parliament in March 2024 (European Parliament, 2024b) and amended as part of the so-called corrigendum procedure (European Parliament, 2024a). We will therefore use this amended version but will refer to earlier versions when relevant.
@@
The focus on human oversight seems to stem from the background of ensuring human-centric development and deployment of AI-based systems. One of the most prominent concerns regarding the wide-range deployment of AI-based systems seems to be “that technological development and rationalized efficiency will take place at the cost of human agency and safety or rights” (Enqvist, 2023). An obvious way to counter such a techno-centric deployment of AI would be a human-centric development and deployment of AI-based systems. As Enqvist (Enqvist, 2023) notes: “Human centricity, as it has come to be (broadly) understood, does not only reflect that human needs are to be met by new technologies, but also incorporates the aim to safeguard individual rights and increase human well-being.” Indeed, in computer science, a popular area of research is concerned with human-centered AI (HCAI) where efficient collaboration between humans and AI systems is investigated (Bansal et al., 2019; Buçinca et al., 2021; Ehsan et al., 2022; Lai et al., 2023; Schoeffer et al., 2023; Tahaei et al., 2023).
@@
One of the foundational assumptions for the effectiveness of oversight measures is that human oversight should not be an end in itself (Laux, 2023), but that the involvement of humans in contexts where AI-based systems affect high-risk decision-making can make things better in some regard. For example, one hope is that humans may be better than AI-based systems at incorporating ethical considerations and social norms into decision-making contexts (Enqvist, 2023). Another hope is that humans might be better at judging single or unusual cases, for instance, in decisions that would affect individual human beings (Longoni et al., 2019; Baum et al., 2022). Additionally, associating the term “effective” with human oversight shows the hope that a human can effectively contribute to a joint decision situation together with an AI-based system, thus enabling decisions better than those of either the human or the system alone (Bansal et al., 2021; Buçinca et al., 2021; Schoeffer et al., 2023). Other hopes include increased safety (Koulu, 2020), improved accuracy (Jones, 2017), or better trustworthiness and more actual trust (Methnani et al., 2021). The most important objective of human oversight recently seems to be the mitigation of risks to fundamental rights, both directly (as one reading of the AI Act (European Parliament, 2024a) suggests), or indirectly via, e.g., the goal to improve safety or the avoidance of ethically and socially undesirable outcomes (Enqvist, 2023).
@@
However, these foundational assumptions have received substantial criticism, calling into question the usefulness of human oversight (Green, 2022; Zerilli et al., 2019). Critics (in particular (Green, 2022)) point to a lack of empirical evidence showing that we, as of now, can reliably integrate human and system abilities in a way that leads to better decisions compared to decisions that every single entity would produce in isolation (Bartlett and McCarley, 2020; Bansal et al., 2021; Rieger and Manzey, 2022). For instance, there seem to be substantial challenges with respect to achieving adequate trust in AI-based systems (Bansal et al., 2019; Lee and See, 2004; Hoff and Bashir, 2015; Zhang et al., 2020). Humans may overly rely on the outputs produced by automated systems, leading to situations where they do not detect erroneous or unfair outputs (Parasuraman and Manzey, 2010). In other situations, humans may have too little trust in AI-based systems or too high confidence in their own abilities, which leads to people overriding actually accurate or fair system outputs (Green and Chen, 2019). Especially in high-reliability contexts, automation bias (cf. Section 4.2) is a likely-to-observe phenomenon where people tend to use the outcome of an automated decision entity “as a heuristic replacement for vigilant information seeking and processing” ((Mosier et al., 1996) p. 205). This makes it particularly challenging for people to detect and adequately address system failures or erroneous outputs. Whereas research has identified several avenues to improve human abilities to oversee systems (e.g., training of oversight persons (Bahner et al., 2008) or augmenting system outputs with confidence scores (Bahner et al., 2008; Bansal et al., 2019)), the positive effects associated with these avenues seem to be inconsistent (Green, 2022). The criticism regarding human oversight could be expressed in an even more foundational way: The driving momentum behind many AI-based systems is the intention of producing decisions and actions that are, in some regard, better than what human experts could achieve (Green, 2022; Zerilli et al., 2019). Ironically, the very human beings whose imperfections are meant to be overcome by such an AI-based system end up in the role of overseers of the system (Zerilli et al., 2019). This is, for instance, a typical situation in certain areas of AI-based medical diagnostics (Haggenmüller et al., 2021; Nishida et al., 2022).
@@
However, it is also possible to draw more positive conclusions from the literature. There are relevant examples where human operators successfully identified errors made by automated systems (McBride et al., 2013; De-Arteaga et al., 2020). In addition, research has enhanced our understanding of factors that affect human abilities to detect errors made by automated systems (De-Arteaga et al., 2020; McBride et al., 2013; Lai et al., 2023; Schoeffer et al., 2023). Furthermore, ongoing research contributes to successfully unveiling design options to improve human-system collaboration (Lai et al., 2023; Vasconcelos et al., 2023). For example, recent work proposes to combine the strengths of machine and human decision-making by using so-called conformal predictions for multi-label classifiers. There, the machine proposes a set of outputs (instead of a single output), which is guaranteed to contain the correct output with a probability of at least some pre-defined value (Straitouri et al., 2023). If this set contains only a single output, confidence is considered high enough to let the decision happen fully autonomously. If, instead, it contains more than one output, the choice is delegated to a human. It was shown that for certain tasks, the human-system decision accuracy indeed is higher when the human choice is restricted to the outputs that are in the proposed set (Straitouri and Rodriguez, 2023).
Overall, research reports mixed results regarding human oversight. However, the fact that effective human oversight is a popular requirement in legislation emphasizes the necessity of obtaining clarity on the very concept of effective human oversight. This clarity is crucial to better devise powerful laws and guidelines and for ensuring the successful and safe implementation of AI in high-risk contexts. Thus, in this paper, we aim to derive high-level conditions for effectiveness in human oversight.
Moral Responsibility
While the approach to effectiveness in human oversight that we suggest in this paper is the original result of collective insights of an interdisciplinary team of researchers from psychology, law, philosophy, and computer science, we also draw inspiration from several high-level conceptions of moral responsibility, as, e.g., in (Talbert, 2019; Noorman, 2020; Fischer and Ravizza, 1998). Most conceptions of moral responsibility have a focus on three aspects in common: a causal connection between the agent and the event for which the agent might be responsible, the agent having sufficient knowledge of their decision situation, and a certain autonomy of the agent (Talbert, 2019; Noorman, 2020; Fischer and Ravizza, 1998). This high-level conception is used as a rough guardrail in this paper.
@@ 3. What is Effective Human Oversight?
We understand human oversight to be the supervision of a system by at least one natural person, typically with the authority to influence its operations or effects. [^2] While this understanding of human oversight will serve as our working definition, it is not the only possible candidate (cf. Section 6). For further analysis of the concept of human oversight, see also (Laux, 2023). This influence operates at various levels throughout the system’s lifecycle, including intervention during execution, reversal of faulty decisions, or adjustments of parameters to improve results (Methnani et al., 2021). Although the oversight person may be assigned additional responsibilities, such as monitoring system performance or making engineering decisions, these tasks are not themselves part of human oversight. Even when having fixated a meaning for “human oversight”, it is not clear what “effective” means in this regard.
3.1. An objective-first approach to the effectiveness of human oversight
@@
We want oversight to be useful. Therefore, it makes sense to define the effectiveness of human oversight in terms of how well it achieves its objectives. As we have discussed above, risk mitigation can be taken to be one of the main objectives of human oversight. This is the objective we will be focusing on. Additionally, effective human oversight arguably may also yield other positive effects, such as improved incorporation of ethical considerations (Enqvist, 2023), or enhanced judgment in cases impacting individual well-being (Longoni et al., 2019; Baum et al., 2022).
@@
It is important to note that the understanding of effectiveness we are about to suggest is not the only viable one. While offering a structured framework, it is just one among potentially numerous conceptualizations. One way to change our conceptualization of effectiveness in human oversight would be to put objectives other than risk mitigation into focus, such as improved accuracy, trustworthiness, actual trust, human autonomy, accountability and responsibility, or liability. Discussing them is beyond the scope of this paper; however, prima facie, there appears to be no inherent conflict between our proposal and these objectives. Furthermore, we believe that our conceptualization will be helpful beyond risk mitigation since it will also be able to address other objectives of human oversight, including (moral) responsibility, as we will demonstrate in Section 3.2.2.
@@
In line with the proposition that the main objective of effective human oversight is to mitigate risk, the main question is: When does human oversight facilitate the mitigation of risks? We propose this is the case if the oversight person meets the following four key conditions: causal power, epistemic access, self-control, and fitting intentions. We believe that this makes intuitive sense: Usually, risks can be reliably mitigated by a human only if there is an action available to them that averts the risk (causal power), they can actually perform that action (self-control), they know that the risk is imminent and which action could mitigate it (epistemic access), and they also want to mitigate it (fitting intentions). This intuition can be put more precisely, as is done in the remainder of this subsection. Additionally, Figure 1 displays the four conditions of effectiveness.
[…]
A diagram that shows the four conditions of effectiveness in human oversight: the boxes for epistemic access, self-control, causal power have outgoing arrows that jointly point to another box for moral responsibility; the box for moral responsibility and the box for fitting intentions have outgoing arrows that jointly point to another box for effectiveness in human oversight.
@@
Figure 1. The four conditions of effectiveness in human oversight and their relation to moral responsibility. The person in the diagram depicts the human in charge of oversight and the graph a schematic depiction of their current decision situation: the node on the left is their current position, the edges are the actions that they could perform, and the nodes on the right are the consequences of their actions.